OpenStudy (anonymous):

I really need help, I am supposed to finish this class soon and im no where near finished. Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively. @lexlover123

OpenStudy (anonymous):

@wintry

OpenStudy (amistre64):

you should have a standard formula that relates to a geometric sequences

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

@Stiwan

OpenStudy (anonymous):

well, there's the following formula for geometric sequences: \[a_n = a_1 \cdot q^{n-1}\]If you want to know the thoughts behind this, just ask. Anyway, now we know \[a_2 = a_1 \cdot q = -21 \]\[a_5 = a_1 \cdot q^4 = 567 \] So we've got 2 equations and 2 unknowns. Can you solve this on your own?

OpenStudy (anonymous):

yes thank you so much @Stiwan